OpenStudy (anonymous):

If f(x) = 8 – 3|x + 2|, find f(–6). If you help me then i will fan/medal.

OpenStudy (anonymous):

A. –16 B. –4 C. 4 D. 20

Directrix (directrix):

f(x) = 8 – 3|x + 2| f(-6) = 8 - 3 * | (-6) + 2| f ( -6) = 8 - 3 * | -4| f(-6) = ? @doresmith

OpenStudy (anonymous):

is it a?

Directrix (directrix):

Please work this out and we will know the answer from your work, okay? f ( -6) = 8 - 3 * | -4|

Directrix (directrix):

What does this equal: 8 - 3 * | -4|

OpenStudy (anonymous):

-20

Directrix (directrix):

No. ===== ==> What is the absolute value of 4?

OpenStudy (anonymous):

-4.0

Directrix (directrix):

No. The absolute value of -4 is +4. The absolute value bars strip off the negative sign.

Directrix (directrix):

8 - 3 * 4 = ? What is this equal to?

Directrix (directrix):

Multiply 3 times 4 and subtract what you get from 8 .

Directrix (directrix):

What did you get?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

is that the answer?

Directrix (directrix):

No.

Directrix (directrix):

Multiply 3 times 4 and subtract what you get from 8 . 8 - 12 = ?

OpenStudy (anonymous):

is it B?

Directrix (directrix):

12 minus 8 = 4. 8 minus 12 does not equal 4.

Directrix (directrix):

Again, what is this: 8 - 12 = ?

OpenStudy (anonymous):

it = -4

Directrix (directrix):

Correct.